OpenStudy (anonymous):

system of equations is given below. 2x + 7y = 1 -3x – 4y = 5 A:Create an equivalent system of equations by replacing the first equation by multiplying the first equation by an integer other than 1, and adding it to the second equation. B:Use any method to solve the equivalent system of equations (the new first equation with the original second equation). C:Prove that the solution for the equivalent system is the same as the solution for the original system of equations.

OpenStudy (anonymous):

Choose one variable to eliminate. X in this case would be easier than Y. 6 would be the lowest common multiple of 2 and 3 so cross multiply both equations. New equations 6x + 21y= 3 -6x-8y = 10 since 6x -6x=0 Add equations together to eliminate the x variable 13y=13 y=1 substitute y back into the first equation x=-3 substitute x and y into original equations to see if they give the correct solutions.

OpenStudy (anonymous):

thanks :) @hopelovelift